OpenStudy (anonymous):

Please =simplify 3√1024...???

OpenStudy (callisto):

Is it (i) \(\sqrt[3]{1024}\) or (ii) \(3\sqrt{1024}\) ??

OpenStudy (unklerhaukus):

\[1024=2\times512=2^2\times256=2^3\times128=2^4\times64\]\[\qquad=2^5\times32=2^6\times 16=2^7\times8=2^8\times4\]\[=2^9\times2=2^{10}\]

OpenStudy (unklerhaukus):

\[2^{10}=\left(2^5\right)^2\]

OpenStudy (unklerhaukus):

\[3\sqrt{1024}=3\sqrt{\left(2^5\right)^2}=\dots\]

OpenStudy (anonymous):

Callisto first option

OpenStudy (anonymous):

Can you answer that?

OpenStudy (anonymous):

do you have the answer?

OpenStudy (callisto):

\[\sqrt[3]{1024} = (2^{10})^{\frac{1}{3}}=2\times 2\times 2\times 2^{\frac{1}{3}} = ...?\]

OpenStudy (callisto):

@Stanny Don't just look for the answer to your question. It's against our code of conduct.

OpenStudy (unklerhaukus):

\[\sqrt[3]{1024}=\sqrt[3]{2^{10}}=\sqrt[3]{2^9\times2}=\sqrt[3]{\left(2^3\right)^3\times2}=\]

OpenStudy (anonymous):

10.08

OpenStudy (callisto):

@sarah43 Posting answer without guiding the asker is also against our code of conduct. Please refer to our code of conduct at http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

Thank you People for help

OpenStudy (anonymous):

ok i am sorry.. just see what number cubed will give you 1024 example cube root of 125 is 5 because 5*5*5=125